# Spin-Orbit Misalignments of Eccentric Black Hole Mergers in AGN Disks

### v1

@@ Abstract
The disks of Active Galactic Nuclei (AGNs) provide a natural environment where stellar mass-black holes (BHs) can dynamically pair, undergo repeated interactions, and eventually merge. It is commonly assumed that gas accretion will both efficiently spin-up disk embedded black holes and align the orbits of embedded binaries with the disk plane, leading to mergers with preferentially positive effective spin parameters ($\chi_{\rm eff}$). Such predictions have motivated the use of $\chi_{\rm eff}$ as a diagnostic for identifying candidate AGN-embedded mergers in the LIGO-Virgo-KAGRA gravitational wave catalog. In this work, we perform post-Newtonian N-body simulations of nearly planar binary-single encounters and apply an empirically motivated gas-driven alignment prescription to characterize the expected $\chi_{\rm eff}$-eccentricity correlations of AGN-embedded mergers. By comparing the alignment and gravitational wave inspiral timescales, we identify the regions of parameter space, across both disk location and binary properties, where full disk-spin-orbit alignment is effective and where it is not. We find that quasi-circular binaries typically align by the time they merge, supporting the standard picture of spin-orbit aligned orientations. By contrast, eccentric binaries (with in-band eccentricity $e_{\rm 10Hz}\gtrsim 0.1$) typically inspiral too quickly for gas torques to act, preserving the post-encounter spin–orbit misalignments and yielding more isotropic $\chi_{\rm eff}$ distributions when disk densities and torque efficiencies are modest. This interplay naturally establishes a correlation between binary eccentricity and $\chi_{\rm eff}$ in AGN disks, highlighting a new key observables of the AGN channel and a potential explanation for massive events such as GW190521 and GW231123.
@@
pubyear: 2025pagerange: Spin-Orbit Misalignments of Eccentric Black Hole Mergers in AGN DisksReferences
@@ 1 Introduction
A decade after the first gravitational wave discovery (Abbott et al., 2016), the LIGO/Virgo/KAGRA (LVK) catalog has grown rapidly, revealing a diverse population of binary black hole (BBH) mergers. Yet, the dominant pathway that produces these systems remains uncertain, with different proposed scenarios. Competing theories include isolated binary evolution in galactic fields (Dominik et al., 2012, 2013, 2015; Kinugawa et al., 2014; Belczynski et al., 2016b, a; Silsbee & Tremaine, 2017; Murguia-Berthier et al., 2017; Rodriguez & Antonini, 2018; Schrøder et al., 2018), primordial black holes (e.g. Sasaki et al., 2016; Bird et al., 2016; Clesse & García-Bellido, 2017), dynamical interactions encounters in globular clusters (GCs) (Portegies Zwart & McMillan, 2000; Lee et al., 2010; Banerjee et al., 2010; Tanikawa, 2013; Bae et al., 2014; Rodriguez et al., 2015, 2016a, 2016b, 2016b; Askar et al., 2017; Park et al., 2017; Samsing, 2018; D’Orazio & Samsing, 2018a; Samsing & D’Orazio, 2018; Samsing et al., 2020; Rozner & Perets, 2022), mergers occurring in the nuclear star clusters (NSCs) of Galactic Nuclei (O’Leary et al., 2009; Hong & Lee, 2015; VanLandingham et al., 2016; Antonini & Rasio, 2016; Stephan et al., 2016; Hoang et al., 2018; Hamers et al., 2018) and in the disks of Active Galactic Nuclei (AGNs) (Bartos et al., 2017; Stone et al., 2017; McKernan et al., 2018; Tagawa et al., 2020a; Samsing et al., 2022; Rozner et al., 2023; Li et al., 2023; DeLaurentiis et al., 2023; Whitehead et al., 2024; Rozner et al., 2023; Rowan et al., 2023, 2024a, 2024b, 2025b; Fabj & Samsing, 2024; Wang et al., 2025b, a; Delfavero et al., 2025b). Each channel is expected to leave characteristic imprints on the properties of the binary that in theory can be used to distinguish between dominant merger origins. The primary properties of relevance are the merger’s mass ratios, orbital eccentricity, and spin orientations. Field binaries formed in isolation are expected to mainly produce approximately equal-mass systems merging on circular orbits, due to their evolution through mass transfer and common envelope phases (e.g. Dominik et al., 2012). In addition, their individual spin components are expected to be predominantly aligned with the direction of the orbital angular momentum (e.g. Kalogera, 2000; Tagawa et al., 2020b; Vajpeyi et al., 2022), resulting in a distribution of positive values for the effective spin parameter $\chi_{\rm eff}$. On the other hand, binaries formed through dynamical interactions and strong few-body encounters inside dynamical environments such as GCs and NSCs, are expected to merge with a variety of mass ratios and retain significant eccentricity in the LVK observable band (e.g. Cholis et al., 2016; O’Leary et al., 2009; Samsing et al., 2014; Samsing, 2018). In addition, dynamical mergers are expected to have an isotropic distribution of spin orientations, producing $\chi_{\rm eff}$ values clustered around zero with both positive and negative tails (e.g. Rodriguez et al., 2016c). However, the analysis of detected sources indicates that neither isolated binary evolution nor dynamical formation within globular or nuclear star clusters alone can fully explain the characteristics of the black hole population observed so far (Zevin et al., 2021). This discrepancy motivates the search for additional pathways capable of reproducing the observed diversity. One promising example is the AGN channel, which combines aspects of both isolated and dynamical evolution within a gas-rich environment.
@@
AGN disks represent a more complex, and yet very promising, scenario that combines elements of both the isolated and dynamical channel of binary formation and merger, with the addition of environmental effects due to e.g. the gaseous disk (Li et al., 2023; DeLaurentiis et al., 2023; Rozner et al., 2023; Rowan et al., 2023, 2024b; Whitehead et al., 2024, 2025b) and tides from the central supermassive black hole (SMBH) (Trani et al., 2019, 2024; Rom et al., 2024; Fabj & Samsing, 2024). In an active galaxy, black holes in the NSC will be captured into the disk as a result of gas-drag processes (Fabj et al., 2020; Nasim et al., 2023; Generozov & Perets, 2023; Wang et al., 2024; Rowan et al., 2025a; Whitehead et al., 2025a). Once fully embedded, they can migrate towards the SMBH, grow to large masses, and undergo repeated encounters (scatterings) potentially yielding multi-generation and hierarchical mergers.
@@
Orbits of stars residing in the NSC can also be aligned with the plane of the disk and undergo modified stellar evolution due to gas accretion (Jermyn et al., 2021; Cantiello et al., 2021; Dittmann et al., 2021; Jermyn et al., 2022; Dittmann et al., 2023; Fabj et al., 2025; Dittmann & Cantiello, 2025), producing a large fraction of massive stars and black holes. Gas not only tends to align inclined stellar objects but it will also align spin-orbit tilted binaries (e.g. Tanaka & Ward, 2004; Tagawa et al., 2020b; McKernan & Ford, 2024), which can originate from dynamical interactions. This implies that a population of AGN-channel mergers would be biased toward a positive skewness for $\chi_{\rm eff}$ (e.g. Tagawa et al., 2020a; McKernan et al., 2024; Cook et al., 2024).
@@
Two events stand out as possible AGN-channel sources. The first is GW190521, one of the most massive BBH mergers detected to date, with a total mass of about $150\,\rm M_{\odot}$ (Abbott et al., 2020). Its inferred properties represent a challenge to the standard isolated binary and dynamical formation scenarios because the component masses lie in or above the pair-instability mass gap (e.g. Fishbach & Holz, 2020; Fishbach et al., 2020; Nitz & Capano, 2021), the waveform is consistent with a possible eccentric merger of $e\geq 0.1$ at 10 Hz (e.g. Romero-Shaw et al., 2020, 2022; Gayathri et al., 2022), and the measured spin–orbit tilt suggests a large misalignment between the component spins and the orbital plane (Abbott et al., 2020). The combination of features have motivated the idea that GW190521 could have formed in an AGN disk, where hierarchical mergers (Tagawa et al., 2021) and accretion (McKernan et al., 2012; Gilbaum & Stone, 2022) can grow massive black holes, while few-body encounters can generate eccentric orbits (e.g. Samsing et al., 2022; Fabj & Samsing, 2024; Rowan et al., 2025b).
@@
The second source of interest is GW231123, the most massive BBH merger up to date, $\sim 200\,\rm M_{\odot}$, with evidence for individual highly spinning components (The LIGO Scientific Collaboration et al., 2025). Such rapid spins are difficult to explain in purely dynamical environments, where black holes typically retain low natal spins or experience spin-down during hierarchical mergers (e.g. Gerosa & Fishbach, 2021; Rodriguez et al., 2019), but could naturally arise in AGN disks if sustained accretion efficiently spins up black holes (Delfavero et al., 2025a; Bartos & Haiman, 2025).
@@
Furthermore, the geometry of dynamical interactions in AGN disks plays an essential role in producing a high number of mergers. In the coplanar 2D-disk limit, in contrast with 3D clusters, all objects share the same orbital plane, maximizing angular momentum exchange, the efficiency of hardening, and the production of eccentric mergers (Samsing et al., 2022). At the same time, introducing even a small inclination angle shifts the dynamics toward the three-dimensional scattering regime, suppressing the probability of eccentric mergers by one to two orders of magnitude once the inclination exceeds about a degree (Samsing et al., 2022; Fabj & Samsing, 2024). The occurrence of non-coplanar encounters is quite plausible, as disk turbulence can inhibit full alignment and maintain small but significant orbital inclinations (e.g. Chen & Lin, 2023; Trani & Di Cintio, 2025). This argument motivates the key question addressed in this work: can dynamically assembled binaries in AGN disks, particularly those that merge with significant eccentricity, retain spin–orbit tilts despite the presence of gas? We provide a first answer to the question, investigating the interplay between gas-driven alignment and gravitational wave inspiral while focusing on how this competition shapes the final inclination and spin properties of merging binaries.
@@ 2 Models
First, we describe our choice of AGN disk model, for which we adopt the dense, marginally stable Sirko & Goodman (2003) (hereafter SG) disk as a reference framework for evaluating disk-dependent quantities. We then present our post-Newtonian N-body simulations of three-body scatterings, designed to capture how binary–single encounters inside the disk generate spin–orbit tilts. Next, we introduce the schematic picture of spin–orbit tilt evolution for dynamically assembled binaries in disks, emphasizing the interplay between misalignment through scattering and subsequent re-alignment through gas accretion. Finally, we outline our alignment prescription, in which the angular momentum evolution of the binary is modeled through gas accretion in different disk regimes, allowing us to quantify how efficiently binaries can re-align compared to their gravitational wave inspiral timescales.
@@ 2.1 AGN Disk Model
We consider the SG disk model for a range of SMBH masses ($10^{6}$$10^{9}\,\rm M_{\odot}$), computed with the pAGN package (Gangardt et al., 2024). The SG model extends the classical Shakura & Sunyaev (1973) thin-disk framework to parsec scales while preventing fragmentation by enforcing marginal stability ($Q\gtrsim 1$) through additional heating. This prescription produces disks that are hotter and denser at large radii than standard $\alpha$-disks, with opacity transitions accounted for using tabulated opacity tables (Semenov et al., 2003; Badnell et al., 2005). Compared to other prescriptions, such as the Thompson et al. (2005) model, the SG disks are systematically denser, particularly in their inner regions, between $10^{2}-10^{4}\,\rm R_{s}$, where $R_{s}$ is the Schwarzschild radius. This makes SG a natural choice to highlight the regime in which alignment can take place most efficiently when merging binaries are embedded in the densest parts of AGN disks. As standard parameters in pAGN, we set the viscosity to $\alpha=0.01$ and the Eddington fraction to $l_{\rm E}=0.5$.
@@ 2.2 Three-Body Scatterings
We carry out post-Newtonian (PN) N-body simulations of black hole (BH) binary–single interactions (three-body scatterings) adopting the same methods introduced in Fabj & Samsing (2024) (referred to FS24 from here on). The scattering experiments account for the tidal field of the central SMBH in the equations of motion but neglect the influence of gas during the encounter. The contribution from the SMBH implemented in FS24 extends the analysis of Samsing et al. (2022) on the probability of dynamically forming an eccentric merger based on the constrained geometry in AGN disks. Besides FS24, the overall code has also been extensively tested in previous studies (Samsing et al., 2014, 2017; Samsing, 2018; Samsing et al., 2022), and we refer the reader to Sec. 3 of FS24 for further details on the numerical implementation and simulation setup.
As the specific scope of this paper is to study the binary spin–orbit tilt distribution in a disk setup, we focus on a three-dimensional encounter geometry. The single incoming BH approaches at a small inclination angle a binary located in the disk mid-plane ($\varphi=0.1^{\circ}$, see first panel of Fig. 1). This represents a likely scenario in AGNs, where the three BHs are initially embedded in the accretion disk, and the vertical offset of the incoming object does not exceed the disk opening angle, set by its characteristic vertical thickness. While our chosen $\varphi$ is small, such out-of-plane perturbations are particularly relevant, as even a low inclination can break the symmetry of coplanar scatterings. In fact, the likelihood of producing eccentric mergers is highly sensitive to $\varphi$, with rates dropping by nearly an order of magnitude for $\varphi=1^{\circ}$ compared to the coplanar case (Samsing et al., 2022; Trani et al., 2024; Fabj & Samsing, 2024). Recent hydrodynamical simulations (e.g Whitehead et al., 2025a; Rowan et al., 2025a) have shown that the single objects on lowly inclined orbits ($\varphi<3^{\circ}$) rapidly embed themselves in the AGN disk, within order $\sim$10 orbits around the SMBH, making the assumption of a relatively low $\varphi$ justified. We model the interaction such that the inclined single BH is placed at the Hill radius of the triple system with respect to the SMBH, with a velocity vector equivalent to the coplanar case, but with zero velocity in the $z$ direction at the start of the simulation. We adopt an equal-mass three-body system with component masses of $15M_{\odot}$, and with mass ratios relative to the SMBH in the range $10^{-5}$$10^{-8}$. The binary is placed in the disk mid-plane and is initially circular with a semi-major axis of 1 AU.

### v2

@@ Abstract
The disks of Active Galactic Nuclei (AGNs) provide a natural environment where stellar mass-black holes (BHs) can dynamically pair, undergo repeated interactions, and eventually merge. It is commonly assumed that gas accretion will both efficiently spin-up disk-embedded black holes and align the orbits of embedded binaries with the disk plane, leading to mergers with preferentially positive effective spin parameters ($\chi_{\rm eff}$). Such predictions have motivated the use of $\chi_{\rm eff}$ as a diagnostic for identifying candidate AGN-embedded mergers in the LIGO-Virgo-KAGRA gravitational wave catalog. In this work, we perform post-Newtonian N-body simulations of nearly planar binary-single encounters and apply an empirically motivated gas-driven alignment prescription to characterize the expected $\chi_{\rm eff}$-eccentricity correlations of AGN-embedded mergers. By comparing the alignment and gravitational wave inspiral timescales, we identify the regions of parameter space, across both disk location and binary properties, where full disk-spin-orbit alignment is effective and where it is not. We find that quasi-circular binaries typically align by the time they merge, supporting the standard picture of spin-orbit aligned orientations. By contrast, eccentric binaries (with in-band eccentricity $e_{\rm 10Hz}\gtrsim 0.1$) typically inspiral too quickly for gas torques to act, preserving the post-encounter spin–orbit misalignments and yielding more isotropic $\chi_{\rm eff}$ distributions when disk densities and torque efficiencies are modest. This interplay naturally establishes a correlation between binary eccentricity and $\chi_{\rm eff}$ in AGN disks, highlighting a new key observable of the AGN channel and a potential explanation for massive events such as GW190521 and GW231123.
@@
pubyear: 2025pagerange: Spin-Orbit Misalignments of Eccentric Black Hole Mergers in AGN Disks7
@@ 1 Introduction
A decade after the first gravitational wave discovery (Abbott et al., 2016), the LIGO/Virgo/KAGRA (LVK) catalog has grown rapidly, revealing a diverse population of binary black hole (BBH) mergers. Yet, the dominant pathway that produces these systems remains uncertain, with different proposed scenarios. Competing theories include isolated binary evolution in galactic fields (Dominik et al., 2012, 2013, 2015; Kinugawa et al., 2014; Belczynski et al., 2016b, a; Silsbee and Tremaine, 2017; Murguia-Berthier et al., 2017; Rodriguez and Antonini, 2018; Schrøder et al., 2018), primordial black holes (e.g. Sasaki et al., 2016; Bird et al., 2016; Clesse and García-Bellido, 2017), dynamical interactions encounters in globular clusters (GCs) (Portegies Zwart and McMillan, 2000; Lee et al., 2010; Banerjee et al., 2010; Tanikawa, 2013; Bae et al., 2014; Rodriguez et al., 2015, 2016a, 2016b, 2016b; Askar et al., 2017; Park et al., 2017; Samsing, 2018; D’Orazio and Samsing, 2018a; Samsing and D’Orazio, 2018; Samsing et al., 2020; Rozner and Perets, 2022), mergers occurring in the nuclear star clusters (NSCs) of Galactic Nuclei (O’Leary et al., 2009; Hong and Lee, 2015; VanLandingham et al., 2016; Antonini and Rasio, 2016; Stephan et al., 2016; Hoang et al., 2018; Hamers et al., 2018) and in the disks of Active Galactic Nuclei (AGNs) (Bartos et al., 2017; Stone et al., 2017; McKernan et al., 2018; Tagawa et al., 2020b; Samsing et al., 2022; Rozner et al., 2023; Li et al., 2023; DeLaurentiis et al., 2023; Whitehead et al., 2024; Rozner et al., 2023; Rowan et al., 2023, 2024b, 2024a, 2025b; Fabj and Samsing, 2024; Wang et al., 2025a, b; Delfavero et al., 2025b). Each channel is expected to leave characteristic imprints on the properties of the binary that in theory can be used to distinguish between dominant merger origins. The primary properties of relevance are the merger’s mass ratios, orbital eccentricity, and spin orientations. Field binaries formed in isolation are expected to mainly produce approximately equal-mass systems merging on circular orbits, due to their evolution through mass transfer and common envelope phases (e.g. Dominik et al., 2012). In addition, their individual spin components are expected to be predominantly aligned with the direction of the orbital angular momentum (e.g. Kalogera, 2000; Tagawa et al., 2020a; Vajpeyi et al., 2022), resulting in a distribution of positive values for the effective spin parameter $\chi_{\rm eff}$. On the other hand, binaries formed through dynamical interactions and strong few-body encounters inside dynamical environments such as GCs and NSCs, are expected to merge with a variety of mass ratios and retain significant eccentricity in the LVK observable band (e.g. Cholis et al., 2016; O’Leary et al., 2009; Samsing et al., 2014; Samsing, 2018). In addition, dynamical mergers are expected to have an isotropic distribution of spin orientations, producing $\chi_{\rm eff}$ values clustered around zero with both positive and negative tails (e.g. Rodriguez et al., 2016c). However, the analysis of detected sources indicates that neither isolated binary evolution nor dynamical formation within globular or nuclear star clusters alone can fully explain the characteristics of the black hole population observed so far (Zevin et al., 2021). This discrepancy motivates the search for additional pathways capable of reproducing the observed diversity. One promising example is the AGN channel, which combines aspects of both isolated and dynamical evolution within a gas-rich environment.
@@
AGN disks represent a more complex, and yet very promising, scenario that combines elements of both the isolated and dynamical channel of binary formation and merger, with the addition of environmental effects due to e.g. the gaseous disk (Li et al., 2023; DeLaurentiis et al., 2023; Rozner et al., 2023; Rowan et al., 2023, 2024a; Whitehead et al., 2024, 2025a) and tides from the central supermassive black hole (SMBH) (Trani et al., 2019, 2024; Rom et al., 2024; Fabj and Samsing, 2024). In an active galaxy, black holes in the NSC will be captured into the disk as a result of gas-drag processes (Fabj et al., 2020; Nasim et al., 2023; Generozov and Perets, 2023; Wang et al., 2024; Rowan et al., 2025a; Whitehead et al., 2025b). Once fully embedded, they can migrate towards the SMBH, grow to large masses, and undergo repeated encounters (scatterings) potentially yielding multi-generation and hierarchical mergers.
@@
Orbits of stars residing in the NSC can also be aligned with the plane of the disk and undergo modified stellar evolution due to gas accretion (Jermyn et al., 2021; Cantiello et al., 2021; Dittmann et al., 2021; Jermyn et al., 2022; Dittmann et al., 2023; Fabj et al., 2025; Dittmann and Cantiello, 2025), producing a large fraction of massive stars and black holes. Gas not only tends to align inclined stellar objects but it will also align spin-orbit tilted binaries (e.g. Tanaka and Ward, 2004; Tagawa et al., 2020a; McKernan and Ford, 2024), which can originate from dynamical interactions. This implies that a population of AGN-channel mergers would be biased toward positive values of $\chi_{\rm eff}$ (e.g. Tagawa et al., 2020b; McKernan et al., 2024; Cook et al., 2024).
@@
Two events stand out as possible AGN-channel sources. The first is GW190521, one of the most massive BBH mergers detected to date, with a total mass of about $150\,\rm M_{\odot}$ (Abbott et al., 2020). Its inferred properties represent a challenge to the standard isolated binary and dynamical formation scenarios because the component masses lie in or above the pair-instability mass gap (e.g. Fishbach and Holz, 2020; Fishbach et al., 2020; Nitz and Capano, 2021), the waveform is consistent with a possible eccentric merger of $e\geq 0.1$ at 10 Hz (e.g. Romero-Shaw et al., 2020, 2022; Gayathri et al., 2022), and the measured spin–orbit tilt suggests a large misalignment between the component spins and the orbital plane (Abbott et al., 2020). The combination of features have motivated the idea that GW190521 could have formed in an AGN disk, where hierarchical mergers (Tagawa et al., 2021) and accretion (McKernan et al., 2012; Gilbaum and Stone, 2022) can grow massive black holes, while few-body encounters can generate eccentric orbits (e.g. Samsing et al., 2022; Fabj and Samsing, 2024; Rowan et al., 2025b).
@@
The second source of interest is GW231123, the most massive BBH merger up to date, $\sim 200\,\rm M_{\odot}$, with evidence for individual highly spinning components (The LIGO Scientific Collaboration et al., 2025). Such rapid spins are difficult to explain in purely dynamical environments, where black holes typically retain low natal spins or experience spin-down during hierarchical mergers (e.g. Gerosa and Fishbach, 2021; Rodriguez et al., 2019), but could naturally arise in AGN disks if sustained accretion efficiently spins up black holes (Delfavero et al., 2025a; Bartos and Haiman, 2025).
@@
Furthermore, the geometry of dynamical interactions in AGN disks plays an essential role in producing a high number of mergers. In the coplanar 2D-disk limit, in contrast with 3D clusters, all objects share the same orbital plane, maximizing angular momentum exchange, the efficiency of hardening, and the production of eccentric mergers (Samsing et al., 2022). At the same time, introducing even a small inclination angle shifts the dynamics toward the three-dimensional scattering regime, suppressing the probability of eccentric mergers by one to two orders of magnitude once the inclination exceeds about a degree (Samsing et al., 2022; Fabj and Samsing, 2024). The occurrence of non-coplanar encounters is quite plausible, as disk turbulence can inhibit full alignment and maintain small but significant orbital inclinations (e.g. Chen and Lin, 2023; Trani and Di Cintio, 2025). This argument motivates the key question addressed in this work: can dynamically assembled binaries in AGN disks, particularly those that merge with significant eccentricity, retain spin–orbit tilts despite the presence of gas? We provide a first answer to the question, investigating the interplay between gas-driven alignment and gravitational wave inspiral while focusing on how this competition shapes the final inclination and spin properties of merging binaries.
@@ 2 Models
First, we describe our choice of AGN disk model, for which we adopt the dense, marginally stable Sirko and Goodman (2003) (hereafter SG) disk as a reference framework for evaluating disk-dependent quantities. We then present our post-Newtonian N-body simulations of three-body scatterings, designed to capture how binary–single encounters inside the disk generate spin–orbit tilts. Next, we introduce the schematic picture of spin–orbit tilt evolution for dynamically assembled binaries in disks, emphasizing the interplay between misalignment through scattering and subsequent re-alignment through gas accretion. Finally, we outline our alignment prescription, in which the angular momentum evolution of the binary is modeled through gas accretion in different disk regimes, allowing us to quantify how efficiently binaries can re-align compared to their gravitational wave inspiral timescales.
@@ 2.1 AGN Disk Model
We consider the SG disk model for a range of SMBH masses ($10^{6}$$10^{9}\,\rm M_{\odot}$), computed with the pAGN package (Gangardt et al., 2024). The SG model extends the classical Shakura and Sunyaev (1973) thin-disk framework to parsec scales while preventing fragmentation by enforcing marginal stability ($Q\gtrsim 1$) through additional heating. This prescription produces disks that are hotter and denser at large radii than standard $\alpha$-disks, with opacity transitions accounted for using tabulated opacity tables (Semenov et al., 2003; Badnell et al., 2005). Compared to other prescriptions, such as the Thompson et al. (2005) model, the SG disks are systematically denser, particularly in their inner regions, between $10^{2}-10^{4}\,\rm R_{s}$, where $R_{s}$ is the Schwarzschild radius. This makes SG a natural choice to highlight the regime in which alignment can take place most efficiently when merging binaries are embedded in the densest parts of AGN disks. In environments where embedded compact objects provide significant local heating, the disk may naturally approach a marginally stable configuration as well (e.g. Epstein-Martin et al., 2025), further motivating the focus on the SG prescription. As standard parameters in pAGN, we set the viscosity to $\alpha=0.01$ and the Eddington fraction to $l_{\rm E}=0.5$.
@@ 2.2 Three-Body Scatterings
We carry out post-Newtonian (PN) N-body simulations of black hole (BH) binary–single interactions (three-body scatterings) adopting the same methods introduced in Fabj and Samsing (2024) (referred to FS24 from here on). The scattering experiments account for the tidal field of the central SMBH in the equations of motion but neglect the influence of gas during the encounter. The contribution from the SMBH implemented in FS24 extends the analysis of Samsing et al. (2022) on the probability of dynamically forming an eccentric merger based on the constrained geometry in AGN disks. Besides FS24, the overall code has also been extensively tested in previous studies (Samsing et al., 2014, 2017; Samsing, 2018; Samsing et al., 2022), and we refer the reader to Sec. 3 of FS24 for further details on the numerical implementation and simulation setup.
As the specific scope of this paper is to study the binary spin–orbit tilt distribution in a disk setup, we focus on a three-dimensional encounter geometry. The single incoming BH approaches at a small inclination angle a binary located in the disk mid-plane ($\varphi=0.1^{\circ}$, see first panel of Fig. 1). This represents a likely scenario in AGNs, where the three BHs are initially embedded in the accretion disk, and the vertical offset of the incoming object does not exceed the disk opening angle, set by its characteristic vertical thickness. While our chosen $\varphi$ is small, such out-of-plane perturbations are particularly relevant, as even a low inclination can break the symmetry of coplanar scatterings. In fact, the likelihood of producing eccentric mergers is highly sensitive to $\varphi$, with rates dropping by nearly an order of magnitude for $\varphi=1^{\circ}$ compared to the coplanar case (Samsing et al., 2022; Trani et al., 2024; Fabj and Samsing, 2024). Recent hydrodynamical simulations (e.g Whitehead et al., 2025b; Rowan et al., 2025a) have shown that the single objects on lowly inclined orbits ($\varphi<3^{\circ}$) rapidly embed themselves in the AGN disk, within order $\sim$10 orbits around the SMBH, making the assumption of a relatively low $\varphi$ justified. We model the interaction such that the inclined single BH is placed at the Hill radius of the triple system with respect to the SMBH, with a velocity vector equivalent to the coplanar case, but with zero velocity in the $z$ direction at the start of the simulation. We adopt an equal-mass three-body system with component masses of $15M_{\odot}$, and with mass ratios relative to the SMBH in the range $10^{-5}$$10^{-8}$. The binary is placed in the disk mid-plane and is initially circular with a semi-major axis of 1 AU.